Circular No. 8617

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V477 SCUTI = NOVA SCUTI 2005 No. 2 G. Pojmanski, Warsaw University Astronomical Observatory, reports

his discovery of an apparent nova on CCD images taken on Oct. 11.026 (at V = 12.0) and 13.066 UT (V = 10.4) with a 180-mm-f.l. telephoto lens in the course of the All-Sky Automated Survey (cf. IAUC 8369, 8574), giving the position of the new object as α = 18^{h}38^{m}43^{s}, δ = −12^{o}16′.3 (equinox 2000.0). Pojmanski notes that no object is visible at this position on Digitized Sky Survey images, though the USNO-B1.0 catalogue lists several faint objects (red mag 19 or fainter) around this position; nothing was visible at this position on an ASAS image taken on Oct. 7.055 (limiting mag V = 14). H. Yamaoka, Kyushu University, reports the independent discovery of this apparent nova at mag 10.6 by K. Haseda (Aichi, Japan, 400-mm-f.l. f/4 telephoto lens + TMax 400 film) on photographs taken on Oct. 13.419 and 13.420, the position for the nova given as α = 18^{h}38^{m}44^{s}, δ = −12^{o}16′25′′. T. Puckett, Ellijay, GA, reports that the nova appeared at mag 10.2 on an unfiltered image taken on Oct. 15.99 UT with his 0.60-m reflector, and he provides the following precise position for the nova: α = 18^{h}38^{m}42^{s}.93, δ = −12^{o}16′15′′.6. Visual magnitude estimates: Oct. 14.770, 11.0 (K. Hornoch, Lelekovice, Czech Republic); 14.846, 11.7 (P. Schmeer, Bischmisheim, Germany); 15.813, 11.7 (Schmeer). N. N. Samus, Institute of Astronomy, Russian Academy of Sciences, informs us that the designation V477 Sct has been given to this nova.

R. K. Das, N. M. Ashok, and D. P. K. Banerjee, Physical Research Laboratory (PRL), Navrangpura, report that their near-infrared JHK spectroscopy (range 1.08–2.35 μm) of V477 Sct, obtained with the Mt. Abu 1.2-m telescope (+ PRL Near-Infrared NICMOS3 Imaging Spectrometer), on Oct. 15.75 UT, are typical of a classical nova and show prominent H I emission lines of the Paschen and Brackett series. Preliminary estimates for the Paβ H I line indicate a FWZI of 6000 km/s or more. O I lines, generally seen in classical-novae spectra, are also seen with the continuum fluorescence 1.3166-μm line being stronger than the Lyβ excited fluorescence O I line at 1.1288 μm.

Yamaoka also writes that a spectrogram of V477 Sct, obtained by M. Fujii (Ibara, Okayama, Japan) on Oct. 16.43 UT, shows a reddish continuum with broad emission lines including Hα (FWHM = 2900 km/s), Hβ, and O I 777.3-nm, suggesting that V477 Sct is a classical nova caught in the early decline phase.

2005 October 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green